# Jet schemes, Newton polygons and continued fractions

### v2

@@ Abstract.
We study jet schemes of Newton non-degenerate plane curve singularities. We identify a subgraph of the graph of jet components and show that it can be constructed from walks on the lattice points in the first quadrant of the Cartesian plane. In particular, we determine all the irreducible components of the jet schemes. Furthermore, we prove that this subgraph encodes the embedded topological type of the curve singularity in the plane. Finally, we introduce a generating series defined in terms of the irreducible components of the jet schemes and their (co-)dimensions, and we prove that this series is rational and explicitly determine its poles.
@@
The second author author was partially supported by Fondecyt project 1221535 and Fondecyt Exploración project 13250049.
@@ 1. Introduction
Their structure has been understood for several classes of singularities, i.e., [14, 15, 9, 17, 4, 20]. A distinctive feature in the study of jet schemes is that, in order to extract information about singularities, one must consider them in a dynamical way: letting $m$ vary and carefully analyzing how the geometry of the jet schemes changes with $m.$ This was one of the motivation of the introduction of the graph of jet components [14, 17]. This is a leveled graph which for a variety $X$ is defined as follows: Its vertices at the level $m$ correspond to the irreducible components of $X_{m}$ and an edge is drawn from a vertex v’ of order $m+1$ to a vertex v’ of order $m$ whenever the component corresponding to $v^{\prime}$ maps into the one corresponding to $v$ under the natural truncation map $X_{m+1}\longrightarrow X_{m}.$ It has been shown, for certain classes of singularities, that the data of this graph (in fact, already of a suitable subgraph) can completely determine the embedded topological type of a singularity [14, 17], and in some cases even its analytical type [4]. Moreover, it is closely related to the problem of resolution of singularities and to the embedded Nash problem [5, 6, 15, 19, 11, 16, 2, 12].
@@
The main purpose of this article is to determine the irreducible components of the jet schemes of a Newton non-degenerate plane curve singularity, and to identify a subgraph of the graph of jet components that encodes the embedded topological type of the singularity. Recall that Newton non-degenerate plane curve singularities are those that can be resolved by a toric morphism whose target is the plane. In the case of an irreducible singularity, we show that this problem is related to the following question: let $p,q$ be two coprime positive integers with $0<p<q.$ Denote by $L_{(p,q)}$ the ray emanating from the origin of the plane in the direction of the point $(p,q),$ i.e., the half-line passing through $(p,q)$ with slope $q/p.$
@@
Let $\alpha\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$. Determine the walk in the plane obtained by the broken line which starts at the point $\alpha+(1,1)$ and in each step it goes towards the halfline $L_{(p,q)}$ until reaching it if possible.
@@
See Figure 1 for a graphical intuition of the problem. We show that this walk (uniquely determined by the problem) can be described by a special type of continued fraction defined via iterated Euclidean divisions. The subgraph of the graph of jet components that we seek is “almost” precisely given by the walks described above (as $\alpha$ varies), where $p$ and $q$ are determined by the Newton polygon of the (irreducible) curve; the integers $p$ and $q$ are closely related to the Puiseux pairs. In the reducible case, the walk is slightly more involved, but we obtain an analogous result. In particular, we determine the irreducible components of all the jet schemes. We prove that the data of the graph of jet components determine the embedded topological type of the curve.
@@
The article is organize as follows:
@@
  • In section 6, we prove the results that will be used as lemmas in proving the main theorems in section 7.
  • Finally, in section 8, we compute an invariant: the generating series associated to the jet schemes over the origin of a Newton non-degenerate plane curve singularity and show it is rational and determine its ”poles”.
@@ 2. Jet Schemes and Arc Spaces
In this section, we introduce the jet schemes and arc space of a variety.
Denote by $k$ an algebraically closed field. Denote by $\mathbf{Var}_{k}$ the category of varieties over $k$. Let $X\rightarrow Spec(k)$ be such a variety. For $m\in\mathbf{N}_{0}$, denote by $\mathcal{X}_{m}:\mathbf{Var}_{k}\rightarrow\mathbf{Set}$ the functor defined by
$\mathcal{X}_{m}(Z)=\mathbf{Var}_{k}[Z\times_{k}Spec(k[t]/\langle t^{m+1}\rangle),X]$
@@
$k[t]/\langle t^{n+1}\rangle\rightarrow k[t]/\langle t^{m+1}\rangle$
These morphisms verify: $\pi_{m,n}\circ\pi_{n,p}=\pi_{m,p}$ for $m\geq n\geq p$. The arc space $X_{\infty}$ of $X$ is the limit object of the jet schemes $X_{m}$ along the truncation morphisms $\pi_{m,n}:X_{m}\rightarrow X_{n}$ for $m\geq n$.
@@
Let’s show how to compute the jet schemes of affine hypersurfaces in $k^{n}=\mathbf{A}^{n}$ using the arc space $\mathbf{A}^{n}_{\infty}$. Set $\mathcal{O}_{k^{n}}=k[\mathbf{x}]$ : $\mathbf{x}=(x_{1},...,x_{n})$ the ring of the affine space $k^{n}$ and let $f\in\mathcal{O}_{k^{n}}$ define a radical ideal. Define $X=\{f=0\}\subset k^{n}$ to be the reduced algebraic set whose ring of functions is
$\mathcal{O}_{X}=\mathcal{O}_{k^{n}}/{\langle f\rangle}$
The ring of the arc space of $k^{n}$ is $\mathcal{O}_{\mathbf{A}^{n}_{\infty}}=k[\mathbf{x}^{(0)},\mathbf{x}^{(1)},\mathbf{x}^{(2)},$...] with variables $\mathbf{x}^{(j)}=(x^{(j)}_{1},...,x^{(j)}_{n})$. Introduce the formal power series:
$x_{i}(t)=\sum_{j=0}^{\infty}x_{i}^{(j)}t^{j}\in\mathcal{O}_{\mathbf{A}^{n}_{\infty}}[[t]]$
Expand $f[x_{1}(t),...,x_{n}(t)]=\sum_{m\geq 0}f^{(m)}t^{m}$. Define now $F_{m}$ to be the algebraic set in $\mathbf{A}^{n}_{\infty}$
$F_{m}=\{f^{(j)}=0\}_{0\leq j\leq m}\subset\mathbf{A}^{n}_{\infty}$
Set $\mathbf{A}^{n}_{m}=\{\mathbf{x}^{(j)}=0\}_{j>m}\subset\mathbf{A}^{n}_{\infty}$. The jet scheme of $X$ at the level $m$ is $X_{m}=\mathbf{A}^{n}_{m}\cap F_{m}\subset\mathbf{A}^{n}_{m}$.
In this study, we take $n=2$, $x=x_{1}$, $y=x_{2}$ and every chosen $f\in\mathcal{O}_{k^{2}}$ is reduced. Set $\mathcal{C}=\{f=0\}$. We are interested in the jet schemes at each level $m$ of $\mathcal{C}$ over the origin of $k^{2}$ ie
$\mathcal{C}^{0}_{m}=\mathbf{A}^{2}_{m}\cap F_{m}\cap\{x^{(0)}=y^{(0)}=0\}\subset\mathbf{A}^{2}_{m}$
That’s why, we will introduce: for $\alpha\in\mathbf{N}^{2}$, $\mathcal{H}^{\alpha}_{\infty}$ to be
$\mathcal{H}^{\alpha}_{\infty}=\{x^{(i)}=y^{(j)}=0:0\leq i<\alpha_{1},0\leq j<\alpha_{2}\}\subset\mathbf{A}^{2}_{\infty}$
Also, set $F^{\alpha}_{m}=\{x^{(\alpha_{1})}\neq 0\}\cap\{y^{(\alpha_{2})}\neq 0\}\cap F_{m}\cap\mathcal{H}^{\alpha}_{\infty}\subset\mathbf{A}^{2}_{\infty}$ and
$\mathcal{H}^{\alpha}_{m}=\mathcal{H}^{\alpha}_{\infty}\cap\mathbf{A}^{2}_{m}\subset\mathbf{A}^{2}_{m}$
Moreover, the contact loci $\text{Cont}^{m}(\mathcal{C})^{0}\subset\mathbf{A}^{2}_{\infty}$ is defined as
$\text{Cont}^{m}(\mathcal{C})^{0}=\{x^{(0)}=y^{(0)}=f^{(0)}=...=f^{(m-1)}=0\}\cap\{f^{(m)}\neq 0\}$
If we embedd $\mathbf{A}^{2}_{m-1}\subset\mathbf{A}^{2}_{\infty}$ naturally for all $m$, then one can write
$\text{Cont}^{m}(\mathcal{C})^{0}=\mathcal{C}^{0}_{m-1}\cap\{f^{(m)}\neq 0\}$
Then, studying the irreducible components of $\mathcal{C}^{0}_{m}$ for every $m$ is equivalent to studying those of $\text{Cont}^{m}(\mathcal{C})^{0}$ for all $m\in\mathbf{N}$.
@@ 3. Special Staircase in the Plane
Let $p,q$ be two coprime positive integers such that $0<p<q.$ Let $L_{(p,q)}$ be the halfline which starts at the origin of the plane and whose slope is $q/p;$ it passes through the point $(p,q)$.
We call elementary steps the unit vectors $(1,0)$ (horizontal) and $(0,1)$ (vertical). A broken line is a finite sequence of such elementary steps. In this section we tackle the following question:
@@
Let $\alpha\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$. What is the walk in the plane obtained by the broken line which starts at the point $\alpha+(1,1)$ and in each step it goes towards the halfline $L_{(p,q)}$ until reaching it if possible ?
Somehow, starting at the point $(1,1)$, the walker is always attracted by the halfline $L_{(p,q)}$ but because of the unit of the steps he is obliged to jump over $L_{(p,q)}$ sometimes before reaching it. Note that the case $0<q<p$ is dual where we replace vertical steps by horizontal ones and vice versa.
@@
Let’s start by observing the example where $p=2$, $q=3$ and $\alpha=(0,0)$: the walk goes along the following points: $(1,1),(1,2),(2,2)$, $(2,3)$ as in Figure 1. This can be encoded by the formula:
$(1,1)+\mathbf{(0,1)+(1,0)+(0,1)}=(2,3)$
@@
Figure 1. The walk mentioned in question 1 for $p=2$, $q=3$ and $\alpha=(0,0)$
@@
To answer Question 1 for $0<p<q$ coprime, we consider $\alpha=(0,0)$ and introduce a form of continued fractions which is given by the following iterated Euclidean divisions :
@@
where in the first equality $d_{1}$ and $r_{1}$ are respectively the quotient and the remainder of the Euclidean division of $q$ by $p;$ recursively, for $k\geq 2,$ $d_{k}$ and $r_{k}$ are respectively the quotient and the remainder of the division of $q+r_{k-1}$ by $p;$ in particular we have $0\leq r_{k}<p.$ If we obtain $r_{i}=0$ for some $i\geq 1,$ we stop and we denote the continued fractions
@@
Here $\mathrm{SC}$ stands for staircase; this name is justified in the theorem below.
@@
The walk which answers Question 1 for $\alpha=(0,0)$ is determined by the following steps: starting from $(1,1),$
For $1\leq i\leq p-1$, $d_{i}$ vertical unit steps followed by one horizontal unit step. Finally, $d_{p}-1$ vertical unit steps.
The walk has the form of a staircase and the $d_{k}$’s determine the heights of the steps. This is why we introduced the notation $\mathrm{SC}.$ Before proving the theorem, we need the following lemma.
Lemma 1.
We have the following:
  1. (1)
    For $k=1,\ldots,p-1,$ $r_{k}>0$ and $r_{p}=0;$ in particular
    $\mathrm{SC}\left(\frac{q}{p}\right)=[[d_{1},\ldots,d_{p}]].$
  2. (2)
    $q=d_{1}+\ldots+d_{p}.$
@@
Summing the first $k$ equations in (3), and then dividing by $p,$ we get
$\frac{kq}{p}=d_{1}+\cdots+d_{k}+\frac{r_{k}}{p}$ (2)
Since $p$ and $q$ are coprime, for $k<p,$ the left member in equation (2) is not an integer, hence $r_{k}\not=0.$
For $k=p,$ the left side in equation (2) is an integer, hence the right side also is; this implies that $r_{p}=0$ and $q=d_{1}+\ldots+d_{p}.$
[…]
Proof.
We need to prove that the walk goes along the following points
$(1,1),(1,2),\ldots,(1,d_{1}+1),(2,d_{1}+1)$
$(2,d_{1}+1),(2,d_{1}+2),\ldots,(2,d_{1}+d_{2}+1),(3,d_{1}+d_{2}+1)$
@@
Since we took $p<q,$ the line $L_{p,q}$ is above the point $(1,1),$ and hence the first step should be vertical. It is sufficient to prove that for $k=1,\ldots,p-1,$ we have :
@@
    The point $(p,d_{1}+\cdots+d_{p-1}+d_{p})$ belong to $L_{p,q}$.
The last assertion follows from the equality $q=d_{1}+\ldots+d_{p}$ in lemma 1. Here the expression the point is strictly below (respectively above) $L_{p,q}$ means that the point belongs to the relative interior of the cone generated by $(1,0)^{t}$ and $(p,q)^{t}$ (respectively by $(p,q)^{t}$ and $(0,1)^{t})$. The first assertion is equivalent to $\det((k,d_{1}+\cdots+d_{k})^{t},(p,q)^{t})>0;$ this latter determinant is equal to
@@
$kq-(q-r_{1})-(q+r_{1}-r_{2})-\cdots-(q+r_{k-1}-r_{k})=r_{k};$
we have simply replaced $pd_{i}$ by $q+r_{i-1}-r_{i},$ see the equations (3). Since by lemma 1 we have $r_{k}>0,$ we obtain the first assertion. Similarly,
$\det((p,q)^{t},(k,d_{1}+\cdots+d_{k}+1)^{t})=p-r_{k}>0$
hence we have the second assertion. The last assertion follows from the fact that $\det((k+1,d_{1}+\cdots+d_{k}+1)^{t},(p,q)^{t})=q-p+r_{k}>0$. ∎
Finally, to answer Question 1 for all $\alpha$, consider the corollary:
@@
Let $0<p<q$ be coprime integers and $\alpha\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$. Then, Question 1 for $(p,q,\alpha)$ and $(p,q,\alpha_{0})$ : $\alpha_{0}=(0,0)$ have the the same sequence of elementary steps.
@@
Write $\alpha=a\cdot(p,q)$ with $a\geq 0$ such that $\alpha+(1,1)=(ap+1,aq+1)$.
$(p,q)=(1,1)+\sum_{i=1}^{p-1}[d_{i}\cdot(0,1)+(1,0)]+(d_{p}-1)(0,1)$
Also, since $\alpha\in L_{(p,q)}$ then $\alpha+(p,q)\in L_{(p,q)}$ and we obtain:
$\alpha+(1,1)+\sum_{i=1}^{p-1}[d_{i}\cdot(0,1)+(1,0)]+(d_{p}-1)(0,1)=\alpha+(p,q)$
Which concludes the proof. ∎

### v3

@@ Abstract.
We study jet schemes of Newton non-degenerate plane curve singularities. We identify a subgraph of the graph of jet components and show that it can be constructed from walks on the lattice points in the first quadrant of the Cartesian plane. In particular, we determine all the irreducible components of the jet schemes. Furthermore, we prove that this subgraph encodes the embedded topological type of the curve singularity in the plane. Finally, we introduce a generating series defined in terms of the irreducible components of the jet schemes and their (co-)dimensions, and we prove that this series is rational and explicitly determines its poles.
@@
The second author author was partially supported by Fondecyt project 1221535 and Fondecyt Exploración project 13250049. The fourth author was partially supported by the ANR-SINTROP
@@ 1. Introduction
Their structure has been understood for several classes of singularities, e.g., [14, 15, 9, 17, 4, 20]. A distinctive feature in the study of jet schemes is that, in order to extract information about singularities, one must consider them in a dynamical way: letting $m$ vary and carefully analyzing how the geometry of the jet schemes changes with $m.$ This was one of the motivation behind introducing the concept known as the graph of jet components [14, 17]. This is a leveled graph which for a variety $X$ is defined as follows: Its vertices at the level $m$ correspond to the irreducible components of $X_{m}$ and an edge is drawn from a vertex $v^{\prime}$ of order $m+1$ to a vertex $v$ of order $m$ whenever the component corresponding to $v^{\prime}$ maps into the one corresponding to $v$ under the natural truncation map $X_{m+1}\longrightarrow X_{m}.$ It has been shown, for certain classes of singularities, that the data of this graph (in fact, already of a suitable subgraph) can completely determine the embedded topological type of a singularity [14, 17], and in some cases even its analytical type [4]. Moreover, it is closely related to the problem of resolution of singularities and to the embedded Nash problem [5, 6, 15, 19, 11, 16, 2, 12].
@@
The main purpose of this article is to determine the irreducible components of the jet schemes of a Newton non-degenerate plane curve singularity, and to identify a subgraph of the graph of jet components that encodes the embedded topological type of the singularity. Recall that Newton non-degenerate plane curve singularities are those that can be resolved by a toric morphism whose target is the plane. In the case of an irreducible singularity, we show that this problem is related to the following problem.
Problem 1.
@@
Let $p,q$ be two coprime positive integers with $0<p<q.$ Denote by $L_{(p,q)}$ the ray emanating from the origin of the plane in the direction of the point $(p,q),$ i.e., the half-line passing through $(p,q)$ with slope $q/p.$ Let $\alpha\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$. Determine the walk in the plane traced by the broken line (i.e. union of line segments) which starts at the point $\alpha+(1,1)$ and, at each step, it moves towards $L_{(p,q)}$ until reaching it.
@@
See Figure 1 for a graphical intuition of the problem. We show that this walk (uniquely determined by the problem) can be described by a special type of continued fraction defined via iterated Euclidean divisions. The subgraph of the graph of jet components that we seek is “almost” precisely given by the walks described above (as $\alpha$ varies), where $p$ and $q$ are determined by the Newton polygon of the (irreducible) curve; the integers $p$ and $q$ are closely related to the Puiseux pairs. In the reducible case, the walk is slightly more involved, but we obtain an analogous result. In particular, we determine the irreducible components of all the jet schemes. In addition, we prove that the data of the graph of jet components determines the embedded topological type of the curve.
@@
The article is organized as follows:
@@
  • In Section 6, we prove some auxiliary results that will be used in showing the main theorems in Section 7.
  • Finally, in Section 8, we compute the following invariant: the generating series associated to the jet schemes over the origin of a Newton non-degenerate plane curve singularity. Then, we show that it is rational and we determine its ”poles” as well.
@@ 2. Jet Schemes and Arc Spaces
In this section, we introduce the jet schemes and arc space of a variety. Denote by $k$ an algebraically closed field. Denote by $\mathbf{Var}_{k}$ the category of varieties over $k$ and Set that of sets. Let $X\rightarrow Spec(k)$ be such a variety. For $m\in\mathbf{N}_{0}$, denote by $\mathcal{X}_{m}:\mathbf{Var}_{k}\rightarrow\mathbf{Set}$, the functor defined by
$\mathcal{X}_{m}(Z)=\mathbf{Var}_{k}[Z\times_{k}Spec(k[t]/\langle t^{m+1}\rangle),X].$
@@
$k[t]/\langle t^{n+1}\rangle\rightarrow k[t]/\langle t^{m+1}\rangle.$
These morphisms verify: $\pi_{m,n}\circ\pi_{n,p}=\pi_{m,p}$ for $m\geq n\geq p$. The arc space $X_{\infty}$ of $X$ is the limit object in $\textbf{Var}_{k}$ of the jet schemes $X_{m}$ along the truncation morphisms $\pi_{m,n}:X_{m}\rightarrow X_{n}$ for $m\geq n$. To illustrate these notions, we take the following example.
@@
In order to find the jet schemes of an affine hypersurfaces in $k^{n}=\mathbf{A}^{n}$ (using the arc space $\mathbf{A}^{n}_{\infty}$), we set $\mathcal{O}_{k^{n}}=k[\mathbf{x}]$ : $\mathbf{x}=(x_{1},...,x_{n})$ the ring of the affine space $k^{n}$ and we let $f\in\mathcal{O}_{k^{n}}$ define a radical ideal. Define $X=\{f=0\}\subset k^{n}$ to be the reduced algebraic set whose ring of functions is
$\mathcal{O}_{X}=\mathcal{O}_{k^{n}}/{\langle f\rangle}.$
The ring of the arc space of $k^{n}$ is $\mathcal{O}_{\mathbf{A}^{n}_{\infty}}=k[\mathbf{x}^{(0)},\mathbf{x}^{(1)},\mathbf{x}^{(2)},$...] with variables $\mathbf{x}^{(j)}=(x^{(j)}_{1},...,x^{(j)}_{n})$. Next, we introduce the formal power series:
$x_{i}(t)=\sum_{j=0}^{\infty}x_{i}^{(j)}t^{j}\in\mathcal{O}_{\mathbf{A}^{n}_{\infty}}[[t]].$
Then, we expand $f[x_{1}(t),...,x_{n}(t)]=\sum_{m\geq 0}f^{(m)}t^{m}$. Now we define $F_{m}$ to be the algebraic set in $\mathbf{A}^{n}_{\infty}$
$F_{m}=\{f^{(j)}=0\}_{0\leq j\leq m}\subset\mathbf{A}^{n}_{\infty}.$
Setting $\mathbf{A}^{n}_{m}=\{\mathbf{x}^{(j)}=0\}_{j>m}\subset\mathbf{A}^{n}_{\infty}$, then the jet scheme of $X$ at the level $m$ is $X_{m}=\mathbf{A}^{n}_{m}\cap F_{m}\subset\mathbf{A}^{n}_{m}$.
In this paper, we take $n=2$, $x=x_{1}$, $y=x_{2}$ and every chosen $f\in\mathcal{O}_{k^{2}}$ is reduced (i.e. with no square factors). We are mainly interested in the jet schemes at each level $m$ of $\mathcal{C}=\{f=0\}$ over the origin of $k^{2}$ i.e.
$\mathcal{C}^{0}_{m}=\mathbf{A}^{2}_{m}\cap F_{m}\cap\{x^{(0)}=y^{(0)}=0\}\subset\mathbf{A}^{2}_{m}.$
For this reason, we shall introduce $\mathcal{H}^{\alpha}_{\infty}$ for $\alpha\in\mathbf{N}^{2}$, to be
$\mathcal{H}^{\alpha}_{\infty}=\{x^{(i)}=y^{(j)}=0:0\leq i<\alpha_{1},0\leq j<\alpha_{2}\}\subset\mathbf{A}^{2}_{\infty}.$
Also, we shall denote $F^{\alpha}_{m}=\{x^{(\alpha_{1})}\neq 0\}\cap\{y^{(\alpha_{2})}\neq 0\}\cap F_{m}\cap\mathcal{H}^{\alpha}_{\infty}\subset\mathbf{A}^{2}_{\infty}$ and
$\mathcal{H}^{\alpha}_{m}=\mathcal{H}^{\alpha}_{\infty}\cap\mathbf{A}^{2}_{m}\subset\mathbf{A}^{2}_{m}.$
Moreover, the contact locus $\text{Cont}^{m}(\mathcal{C})^{0}\subset\mathbf{A}^{2}_{\infty}$ is defined as
$\text{Cont}^{m}(\mathcal{C})^{0}=\{x^{(0)}=y^{(0)}=f^{(0)}=...=f^{(m-1)}=0\}\cap\{f^{(m)}\neq 0\}.$
If we embed $\mathbf{A}^{2}_{m-1}\subset\mathbf{A}^{2}_{\infty}$ naturally for all $m$, then one can write
$\text{Cont}^{m}(\mathcal{C})^{0}=\mathcal{C}^{0}_{m-1}\cap\{f^{(m)}\neq 0\}.$
Thus, studying the irreducible components of $\mathcal{C}^{0}_{m}$ for every $m$ is equivalent to studying those of $\text{Cont}^{m}(\mathcal{C})^{0}$ for all $m\in\mathbf{N}$.
@@ 3. Special Staircase in the Plane
Let $p,q$ be two coprime positive integers such that $0<p<q.$ Let $L_{(p,q)}$ be the half-line which starts at the origin of the plane and whose slope is $q/p;$ it passes through the point $(p,q)$.
We call elementary steps the unit vectors $(1,0)$ (horizontal) and $(0,1)$ (vertical). A broken line is a finite sequence of such elementary steps. In this section, we shall answer the following question:
@@
Let $\alpha\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$. What is the walk in the plane obtained by the broken line which starts at the point $\alpha+(1,1)$ and at each step it moves towards the half-line $L_{(p,q)}$ until reaching it?
It’s worth noting that starting at the point $(1,1)$, the walker is always attracted by the half-line $L_{(p,q)}$. However, because of the unit of the steps he is obliged to jump over $L_{(p,q)}$ sometimes before reaching it. Note that the case $0<q<p$ is symmetric where we replace vertical steps by horizontal ones and vice versa. To clarify our procedures, we take the following examples.
@@
Take $p=2$, $q=3$ and $\alpha=(0,0)$. Then, clearly the walker moves along the following points: $(1,1),(1,2),(2,2)$, $(2,3)$ in that order. This is seen in Figure 1. Clearly, we can encode this fact from the following formula:
$(1,1)+\mathbf{(0,1)+(1,0)+(0,1)}=(2,3).$
@@
Figure 1. The walk mentioned in Question 1 for $p=2$, $q=3$ and $\alpha=(0,0)$.
@@
To answer Question 1 for $0<p<q$ coprime, we start with $\alpha=(0,0)$ and introduce a form of continued fractions which is given by the following iterated Euclidean divisions:
@@
where in the first equality $d_{1}$ and $r_{1}$ are respectively the quotient and the remainder of the Euclidean division of $q$ by $p$. Recursively, for $k\geq 2,$ $d_{k}$ and $r_{k}$ are respectively the quotient and the remainder of the division of $q+r_{k-1}$ by $p$. In particular, we have $0\leq r_{k}<p.$ Once we obtain $r_{i}=0$ for some index $i\geq 1,$ we stop and we denote the continued fractions
@@
Here $\mathrm{SC}$ stands for staircase; this name can be justified in Theorem 1. Before proving the theorem, we need the following lemma.
Lemma 1.
With the same notation as above, we have the following.
For $k=1,\ldots,p-1,$ $r_{k}>0$ and $r_{p}=0$ so that
$\mathrm{SC}\left(\frac{q}{p}\right)=[[d_{1},\ldots,d_{p}]],$
and $q=d_{1}+\ldots+d_{p}.$
Proof.
By summing up the first $k$ equations in (3), and then dividing by $p,$ we get
$\frac{kq}{p}=d_{1}+\cdots+d_{k}+\frac{r_{k}}{p}.$ (2)
Since $p$ and $q$ are coprime, then for $k<p,$ the left-hand side of Equation (2) is not an integer, and hence $r_{k}\not=0.$ However, for $k=p,$ the left-hand side of Equation (2) is an integer so that the right-hand side also is. This implies that $r_{p}=0$ $q=d_{1}+\ldots+d_{p},$ and the proof is complete.
As a result, we have the following theorem.
@@
The walk which answers Question 1 for $\alpha=(0,0)$ is determined by the following description:
  • The walker starts at $(1,1)$,
  • For $1\leq i\leq p-1$, the walker moves vertically $d_{i}$ unit steps followed by one unit step horizontally,
  • Finally, the walker makes $d_{p}-1$ unit steps vertically.
@@
Clearly, it is enough to prove that the walker moves along the following points:
$(1,1),(1,2),\ldots,(1,d_{1}+1),(2,d_{1}+1),$
$(2,d_{1}+2),\ldots,(2,d_{1}+d_{2}+1),(3,d_{1}+d_{2}+1),$
@@
As $p<q$, the line $L_{p,q}$ is above the point $(1,1),$ and hence the first step is vertical. To prove the same for $k=1,\ldots,p-1,$ it is enough to show the following:
@@
    The point $(p,d_{1}+\cdots+d_{p-1}+d_{p})$ belongs to $L_{p,q}$.
Note that the last assertion follows from the equality $q=d_{1}+\ldots+d_{p}$ in Lemma 1. Here the point being strictly below (respectively above) $L_{p,q}$ means that it belongs to the relative interior of the cone generated by $(1,0)^{t}$ and $(p,q)^{t}$ (respectively by $(p,q)^{t}$ and $(0,1)^{t})$ where $t$ stands for transpose. The first assertion is equivalent to $\det\left[\begin{array}[]{cc}k&p\\ d_{1}+\cdots+d_{k}&q\end{array}\right]>0.$ Now by replacing $pd_{i}$ by $q+r_{i-1}-r_{i},$ in (3), this latter determinant is equal to
@@
$kq-(q-r_{1})-(q+r_{1}-r_{2})-\cdots-(q+r_{k-1}-r_{k})=r_{k}.$
From Lemma 1, we know that $r_{k}>0,$ so that we obtain the first assertion. Similarly,
$\det\left[\begin{array}[]{cc}p&k\\ q&d_{1}+\cdots+d_{k}+1\end{array}\right]=p-r_{k}>0$
which gives the second assertion. Finally, the last assertion follows from the fact that $\det\left[\begin{array}[]{cc}k+1&p\\ d_{1}+\cdots+d_{k}+1&q\end{array}\right]=q-p+r_{k}>0$. ∎
Finally, the answer to Question 1 for all $\alpha$ is given in the next corollary.
@@
Let $0<p<q$ be coprime integers and $\alpha\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$. Then, the answer to Question 1 for both $(p,q,\alpha)$ and $(p,q,\alpha_{0})$ with $\alpha_{0}=(0,0)$ is given by the same sequence of elementary steps.
@@
Let $\alpha=a\cdot(p,q)$ with $a\geq 0$, and such that $\alpha+(1,1)=(ap+1,aq+1)$. Clearly,
$(p,q)=(1,1)+\sum_{i=1}^{p-1}[d_{i}\cdot(0,1)+(1,0)]+(d_{p}-1)(0,1).$
As $\alpha\in L_{(p,q)},$ then obviously $\alpha+(p,q)\in L_{(p,q)}.$ Thus,
$\alpha+(1,1)+\sum_{i=1}^{p-1}[d_{i}\cdot(0,1)+(1,0)]+(d_{p}-1)(0,1)=\alpha+(p,q).$
This concludes the proof. ∎